A boy is 10 years older than his brother. In 4 years he will be twice as old as
his brother. Find the present age of each

Answer by stanbon(75887)

A boy is 10 years older than his brother. In 4 years he will be twice as old as
his brother. Find the present age of each
-----
Equation:
y = r + 10
y+4 = 2(r+4)
---------------
Substitute for "y" and solve for "r"::
r+10 + 4 = 2r + 8
------
r = 6 (Brother's age now)
----
y = r+10 = 16 (Boy's age now)
-----------
Cheers,
Stan H.
-------------